OpenStudy (anonymous):

factor the sum or differences of cubes. 1. X^3+1 2. 8x^3-125

OpenStudy (anonymous):

Ok so the rule for sum or difference in cubes is: \[a^3+b^3=(a+b)(a^2-ab+b^2)\] and;\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

So in your first problem what are the two cubes and are they sum or difference??

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok so in your first problem a and b are x and 1 so plug that into the equation for sum of cubes.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Look: \[a^3+b^3=x^3+1^3\]

OpenStudy (anonymous):

\[=(x+1)(x^2-(1)(x)+1^2)\]=\[(x+1)(x^2-x+1)\]

OpenStudy (anonymous):

see how i plugged those into the equation for sum of cubes?

OpenStudy (anonymous):

The way I remember the formula for sum and difference of cubes is using the mnemonic: Same Opposite Always Positive This is referring to the addition or subtraction sign in a^3+b^3 or a^3-b^3: As you can see in the factorization of the addition or sum of cubes the first "operation symbol" is the SAME as the addition symbol (+). The second one is the OPPOSITE of the addition symbol. and the third one is ALWAYS the addition symbol. In the factorization of difference or subtraction of cubes . the first "operation symbol" is the SAME as the subtraction symbol (-). The second one is the OPPOSITE as subtractions ie. (+). and the third one is always POSITIVE ie. (+).

OpenStudy (anonymous):

okay thank you so for the first one the answer would be (x+1)(x^2-x+1)

OpenStudy (anonymous):

Now for your second problem it can be written as:\[(2x)^3+5^3\]so plug that into the difference of cubes formula: \[(2x-5)((2x)^2+(5)(2x)+5^2)\]\[=(2x-5)(4x^2+10x+25)\]

OpenStudy (anonymous):

Yah

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

the SOAP thing kind of?

OpenStudy (anonymous):

why is there a 2 and 5 only

OpenStudy (anonymous):

the question is 8x^3-125

OpenStudy (anonymous):

yes \[8x^3=(2x)^3\]

OpenStudy (anonymous):

law of exponents:\[(ab)^3=a^3b^3\]

OpenStudy (anonymous):

\[(2x)^3=2^3x^3\]

OpenStudy (anonymous):

\[=8x^3\]

OpenStudy (anonymous):

and btw 125 is 5^3

OpenStudy (anonymous):

Just try to memorize the formulas for sum and difference of cubes it will really help you

OpenStudy (anonymous):

i get i now thank you

OpenStudy (anonymous):

glad to help